Circular No. 6942

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

JUPITER J. Lecacheux and P. Drossart, Meudon Observatory; F. Colas, Bureau

des Longitudes; G. Orton and B. Fisher, Jet Propulsion Laboratory; A. Sanchez-Lavega, R. Hueso, and J. F. Rojas, University of Pais Vasco; and I. Miyazaki write: “Jupiter’s 58-year-old white ovals known as BC and DE are now a single white oval located at planetographic latitude −34^{o}. Picdu-Midi observations in the red continuum on 1998 Jan. 17 show a distinct, very close pair of ovals, but infrared observations made with the NASA Infrared Telescope Facility on Mar. 27 (after Jupiter emerged from solar conjunction) show only a single oval. This aspect was confirmed via red and methane-band (890-nm) filters obtained since May 7 at Pic-du-Midi. The new oval ‘BE’ drifts at a rate of −0^{o}.20/day relative to the System III rotation rate of the interior, and the oval’s corresponding longitude can be represented (from data May 7–June 6) by 10338^{o}.2 – 0^{o}.2025 (JD – 2400000). Extrapolation of the new oval position back to the 1997 Dec.–1998 Jan. time frame places it at the location of BC, but further observations are needed the better to constrain the dynamics of the new oval. The collision event took place in Feb. 1998, when Jupiter was close to conjunction. The visible and infrared images show the new oval ‘BE’ larger by ∼ 20 percent than either BC or DE was. White oval FA is unchanged, drifting at a rate of −0^{o}.158/day and separated westward from ‘BE’ by 24^{o} on June 12.”

COMET C/1995 O1 (HALE-BOPP) Ephemeris extension to IAUC 6799 (orbital elements MPC 31204):

1998 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
June 16 6^{h}06.^{m}41 −49^{o}57.^{′}7 5.780 5.578 73.^{o}6 10.^{o}1  9.3
26 6 16.41 −50 25.7 5.870 5.671 73.8  9.9  9.4
July 6 6 26.33 −51 05.6 5.957 5.764 74.2  9.8  9.5
16 6 36.05 −51 56.6 6.041 5.856 74.7  9.6  9.6
26 6 45.46 −52 58.3 6.122 5.947 75.4  9.5  9.7
Aug. 5 6 54.43 −54 09.8 6.200 6.038 76.1  9.4  9.8
15 7 02.82 −55 30.3 6.277 6.128 76.9  9.3  9.9
25 7 10.47 −56 58.7 6.354 6.217 77.7  9.1 10.0
Sept. 4 7 17.18 −58 33.9 6.429 6.306 78.5  9.0 10.0
14 7 22.75 −60 14.4 6.505 6.394 79.3  8.9 10.1
24 7 26.91 −61 58.9 6.580 6.482 80.0  8.8 10.2

1998 June 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green